OpenStudy (sonyalee77):

Medical researchers have determined that for exercise to be beneficial, a person’s desirable heart rate, R, in beats per minute, can be approximated by the formulas R=143-0.65a for women R=165-0.75a for men, where a represents the person’s age. If the desirable heart rate for a man is 135 beats per minute, how old is he? a. 22.5 years old b. 40 years old c. 45 years old d. 42.5 years old Evaluate the function f(x)=1/2x+3 for f(4). a. 5 b. 9 c. 11 d. 14

OpenStudy (sonyalee77):

@texaschic101

OpenStudy (sonyalee77):

@austinL

OpenStudy (sonyalee77):

@Kainui

OpenStudy (kainui):

So you're given the heart rate and you have an equation that relates heart rate to age. So if you plug in heart rate to that equation, you can solve for age.

OpenStudy (texaschic101):

R = 165 - 0.75a (use this formula because we are dealing with a guy) R = 135 now we sub 135 = 165 - 0.75a -- subtract 165 from both sides 135 - 165 = -0.75a - 30 = -0.75a -- now divide by -0.75 -30/-0.75 = a 40 = a

OpenStudy (sonyalee77):

?thanx but can u help me with the other Question

OpenStudy (texaschic101):

f(4) = 1/2(4) + 3 f(4) = 4/2 + 3 f(4) = 2 + 3 = 5

OpenStudy (sonyalee77):

thanx a lot

OpenStudy (texaschic101):

sure thing :)